### v1

@@ IV.4 Double Pulse Signature
Fig. 13 shows the time delays measured between the air and ice pulses as a function of the distance of the antenna to the shower axis for varying antenna depths, using a primary energy $E_{p}=10^{17}$ eV, zenith angle $\theta=0^{\circ}$ and azimuth angle $\phi=0^{\circ}$. To determine the arrival time of either an in-air or in-ice pulse we constructed the Hilbert envelope of the given pulse after applying an 8th order digital Butterworth band-pass filter for the frequency band $30-1000$ MHz. The arrival time was then defined as the earliest time where the envelope reaches $33\%$ of its maximum value.
@@
Figure 13: The difference in arrival time of air and ice radio pulses as a function of distance to the shower axis for varying depths as indicated by the legend, using a primary energy $E_{p}=10^{17}$ eV and a zenith angle $\theta=0$. The arrival time was defined as the time where the Hilbert envelope reaches $33\%$ of its maximum value, after passing an 8th order digital Butterworth band-pass filter for the frequency band $30-1000$ MHz. Circles indicate the arrival time of the North component of the electric field, squares correspond to the West component and triangles correspond to the vertical component.
@@ V Conclusion
Ray tracing was performed analytically by using exponential refractive index profiles for both air and ice, leading to significant ray-bending for in-ice radio propagation. For ice-to-ice ray tracing in general two contributions are found, which we called the direct and indirect emission, the latter consisting of reflected and refracted ray solutions. To achieve reasonable computation times the ray tracing was performed through the interpolation of premade tables, which implies that also slower, more accurate ray tracers can be used.
To account for ray tracing we investigated the behaviour of the end-point formalism, and found that it still holds when correctly interpreting the different variables. We also included the transmission Fresnel coefficients for air-to-ice rays moving through the air-ice boundary, and the reflection Fresnel coefficients for ice-to-ice rays reflecting on the ice-air boundary. Furthermore, we included a focusing factor to account for ray convergence and divergence. Finally, we note that the end-point formalism should naturally lead to the inclusion of transition radiation, as a consequence of how the in-air simulation component and the in-ice simulation component are combined.
We discussed the first simulation results, using a proton with a primary energy of $E_{p}=10^{17}$ eV, a zenith angle $\theta=0^{\circ}$ and an azimuth angle $\phi=0^{\circ}$. The fluence footprint of the in-air emission is marked by the typical bean-shaped pattern arising from the interference between geomagnetic and Askaryan emission, while the in-ice emission fluence footprint shows a symmetric Cherenkov ring associated to Askaryan emission. The combination of both gives a unique footprint showing the bean-shaped pattern from the in-air emission enclosed by the Cherenkov ring from the in-ice emission. We also showed the electric field as a function of time for different polarizations and antenna positions along two different axes, and discussed general features characterizing the in-air and in-ice components. We compared the in-ice radio emission to the result obtained without ray tracing, applied an 8th order digital Butterworth band-pass filter to study the frequency dependence of the emission, and studied the dependence of the different components of the radio emission on the energy of the primary particle. Finally we discussed the typical double pulse signature of the time traces, arising due to the different arrival times of the in-air and in-ice emission. We showed that in general the in-air emission arrives before the in-ice emission, and that the time difference between both pulses shows a clear relation with distance between receiver and shower axis, as well as receiver depth.
FAERIE combines the CORSIKA 7.7500 and the GEANT4 simulation toolkit for a full Monte-Carlo simulation of the radio emission from cosmic-ray particle cascades, taking into account the propagation and emission in both air and ice. The code is available on request. In the future it will be possible to simulate similar cross-media scenarios in one joint step within CORSIKA 8 [57, 58].
@@ VI Acknowledgements
This work was supported by the Flemish Foundation for Scientific Research (FWO-G085820N and I001423N) as well as the European Research Council under the European Unions Horizon 2020 research and innovation program (No 805486 - K. D. de Vries).
@@ Appendix A Shower profiles
[…]
Figures 16 and 17 show the energy deposited in ice by the in-ice particle cascades of respectively the simulated air shower at a primary energy of $E_{p}=10^{17}$ eV and $E_{p}=10^{18}$ eV, both with a zenith angle $\theta=0$.
Figure 14: The number of particles (left) and the distribution of the energy (right) in the air shower as a function of depth with primary energy $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The particle distributions are obtained over the full radial extend of the in-air particle cascade. Similar results were shown in [29].
Figure 15: The number of particles (left) and the distribution of the energy (right) in the air shower as a function of depth with primary energy $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The particle distributions are obtained over the full radial extend of the in-air particle cascade. Similar results were shown in [29].
[…]
@@ Appendix B Atmospheric model
Table 1: The atmospheric density parameters used for the simulations in this work.
[…]
@@
Table 2: The atmospheric index of refraction parameters used for the simulations in this work.
@@ Appendix C CPU time and job run time
Running the simulation of a cosmic-ray cascade using FAERIE requires the three following steps: (1) Running the in-air simulation with CORSIKA and CoREAS, (2) processing the CORSIKA particle output file to generate input files for the GEANT4-based module, and (3) running the in-ice simulation with the GEANT4-based module. Since multiple CPU’s were used during steps (1) and (3) of the simulation process, we can distinguish between CPU time (the total time spent by a single CPU) and the job run time (the time between the start and the end of the simulation process).
To simulate the cosmic-ray cascade initiated by a $E_{p}=10^{17}$ eV proton entering the atmosphere at a zenith angle $\theta=0^{\circ}$ using a star-shaped antenna grid of 121 antennas, the in-air simulation with CORSIKA and CoREAS was split up over 8 different CPU’s. Each CPU simulated the full cosmic-ray air shower and its corresponding radio emission for 15 of the 121 antennas, with one of the CPU’s including a 16th antenna at (0,0). The total CPU time for the in-air simulation was $171$h, which gives average CPU time of $21.3$h. The longest CPU time was $26.8$h, which equals the job run time.
To reduce the CPU time of the in-ice simulation, only the particles within a radius of $1$ m of the shower axis are propagated through the ice [29]. Furthermore we split up the particle footprint into separate parts, running the in-ice simulation for each part on a different CPU. Note that following this approach, a single CPU can simulate the full antenna grid of 121 antennas. The splitting process is based on the energy of the particles, where the amount of particles per part can vary from $540,000$ (energies below $10^{7}$ eV) down to only 1 (starting at energies of $10^{13}$ eV). Figure 18 shows the distribution of the CPU time of the $591$ cores used for the in-ice simulation. Most of the CPU’s manage to stay within $10$ h of CPU time. Some cores however take significantly longer, driving up the job run time to approximately $24$ h. The distribution shows there is still room for optimization in the splitting process. Combining the particle groups corresponding to the simulations with low CPU time would decrease the amount of cores needed, while splitting up the particle groups corresponding to the simulations with high CPU time would decrease the job run time. The longest CPU time for a simulation of the propagation of a single particle was $12.2$ h, which is in principle a hard under limit of the job run time, unless the antenna grid is distributed over different cores as well. The total CPU time was $2985$ h.

### v4

@@ IV.4 Double pulse signature
Figure 13 shows the time delays measured between the air and ice pulses as a function of the distance of the antenna to the shower axis for varying antenna depths, using a primary energy $E_{p}=10^{17}$ eV, zenith angle $\theta=0^{\circ}$ and azimuth angle $\phi=0^{\circ}$. To determine the arrival time of either an in-air or in-ice pulse we constructed the Hilbert envelope of the given pulse after applying an 8th order digital Butterworth bandpass filter for the frequency band $30-1000$ MHz. The arrival time was then defined as the earliest time where the envelope reaches $33\%$ of its maximum value.
@@
Figure 13: The difference in arrival time of air and ice radio pulses as a function of distance to the shower axis for varying depths as indicated by the legend, using a primary energy $E_{p}=10^{17}$ eV and a zenith angle $\theta=0$. The arrival time was defined as the time where the Hilbert envelope reaches $33\%$ of its maximum value, after passing an 8th order digital Butterworth bandpass filter for the frequency band $30-1000$ MHz, averaged over the three components.
@@ V Conclusion
Ray tracing was performed analytically by using exponential refractive index profiles for both air and ice, leading to significant ray bending for in-ice radio propagation. For ice-to-ice ray tracing in general two contributions are found, which we called the direct and indirect emission, the latter consisting of reflected and refracted ray solutions. To achieve reasonable computation times the ray tracing was performed through the interpolation of premade tables, which implies that also slower, more accurate ray tracers can be used.
To account for ray tracing we investigated the behavior of the endpoint formalism, and found that it still holds when correctly interpreting the different variables. We also included the transmission Fresnel coefficients for air-to-ice rays moving through the air-ice boundary, and the reflection Fresnel coefficients for ice-to-ice rays reflecting on the ice-air boundary. Furthermore, we included a focusing factor to account for ray convergence and divergence.
We discussed the first simulation results, using a proton with a primary energy of $E_{p}=10^{17}$ eV, a zenith angle $\theta=0^{\circ}$ and an azimuth angle $\phi=0^{\circ}$. The fluence footprint of the in-air emission is marked by the typical bean-shaped pattern arising from the interference between geomagnetic and Askaryan emission, while the in-ice emission fluence footprint shows a symmetric Cherenkov ring associated to Askaryan emission. The combination of both gives a unique footprint showing the bean-shaped pattern from the in-air emission enclosed by the Cherenkov ring from the in-ice emission. We also showed the electric field as a function of time for different polarizations and antenna positions along two different axes, and discussed general features characterizing the in-air and in-ice components. We compared the in-ice radio emission to the result obtained without ray tracing, applied an 8th order digital Butterworth bandpass filter to study the frequency dependence of the emission, and studied the dependence of the different components of the radio emission on the energy of the primary particle. Finally we discussed the typical double pulse signature of the time traces, arising due to the different arrival times of the in-air and in-ice emission. We showed that in general the in-air emission arrives before the in-ice emission, and that the time difference between both pulses shows a clear relation with distance between receiver and shower axis, as well as receiver depth.
FAERIE combines the CORSIKA 7.7500 and the GEANT4 simulation toolkit for a full Monte-Carlo simulation of the radio emission from cosmic-ray particle cascades, taking into account the propagation and emission in both air and ice. In the future it will be possible to simulate similar cross-media scenarios in one joint step within CORSIKA 8 [61, 62].
@@ VI Acknowledgements
This work was supported by the Flemish Foundation for Scientific Research (FWO-G085820N, FWO-I002119N, and FWO-I001423N) as well as the European Research Council under the European Unions Horizon 2020 research and innovation program (No 805486 - K. D. de Vries).
The code is available on request. A request can be made by contacting the authors via e-mail.
@@ Appendix A Shower profiles
Figures 14 and 15 show the longitudinal particle and energy distributions of respectively the simulated air shower at a primary energy of $E_{p}=10^{17}$ eV and $E_{p}=10^{18}$ eV, both with a zenith angle $\theta=0$. The air-ice boundary is located at a depth of $729$ g/cm2, which means the distributions only show the in-air particle cascade. Figure 16 shows the total energy contained within a given radius from the shower core at the air-ice boundary for both showers. The energy deposited in the ice by both particle cascades is shown in Figure 17.
Figure 14: The number of particles (left) and the distribution of the energy (right) in the air shower as a function of depth with primary energy $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The particle distributions are obtained over the full radial extent of the in-air particle cascade. Similar results were shown in [22].
Figure 15: The number of particles (left) and the distribution of the energy (right) in the air shower as a function of depth with primary energy $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The particle distributions are obtained over the full radial extent of the in-air particle cascade. Similar results were shown in [22].
Figure 16: The total energy contained within a given radius from the shower core at the air-ice boundary up to $100$ m (left) and in more detail up to $5$ m (right), for the simulated shower with a primary energy $E_{p}=10^{17}$ eV and zenith angle $\theta=0$ and the simulated shower with a primary energy $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The energy is expressed as a fraction of the primary energy $E_{p}$. The air-ice boundary is located at an altitude of $2.835$ km, which corresponds to a vertical atmospheric depth of $729$ g/cm2.
Figure 17: The energy deposited in ice by the in-ice particle cascade for the simulated shower with a primary energy $E_{p}=10^{17}$ eV and zenith angle $\theta=0$ (left) and the simulated shower with a primary energy $E_{p}=10^{18}$ eV and zenith angle $\theta=0$ (right). Shown here is the deposited energy density within a vertical 1-cm wide slice going through the center of the particle shower. The air-ice boundary is located at an altitude of $2.835$ km, which corresponds to a vertical atmospheric depth of $729$ g/cm2. Similar results were shown in [22].
@@ Appendix B Atmospheric model
Table 1: The numerical values of the parameters in Equations 10 and 11 describing the density of the atmosphere used for the simulations in this work.
     Layer      Altitude interval (m)      $B_{i}$      $C_{i}$ (m-1)
@@
Table 2: The numerical values of the parameters in Equation 12 describing the refractive index of the atmosphere used for the simulations in this work.
@@ Appendix C CPU time and job run time
Running the simulation of a cosmic-ray cascade using FAERIE requires the three following steps: (1) running the in-air simulation with CORSIKA and CoREAS, (2) processing the CORSIKA particle output file to generate input files for the GEANT4-based module, and (3) running the in-ice simulation with the GEANT4-based module. Since multiple CPUs were used during steps (1) and (3) of the simulation process, we can distinguish between CPU time (the total time spent by a single CPU) and the job run time (the time between the start and the end of the simulation process).
To simulate the cosmic-ray cascade initiated by a $E_{p}=10^{17}$ eV proton entering the atmosphere at a zenith angle $\theta=0^{\circ}$ using a star-shaped antenna grid of 121 antennas, the in-air simulation with CORSIKA and CoREAS was split up over 8 different CPUs. Each CPU simulated the full cosmic-ray air shower and its corresponding radio emission for 15 of the 121 antennas, with one of the CPUs including a 16th antenna at (0,0). The total CPU time for the in-air simulation was $171$h, which gives an average CPU time of $21.3$h. The longest CPU time was $26.8$h, which equals the job run time.
To reduce the CPU time of the in-ice simulation, only the particles within a radius of $1$ m of the shower axis are propagated through the ice [22]. Furthermore we split up the particle footprint into separate parts, running the in-ice simulation for each part on a different CPU. Note that following this approach, a single CPU can simulate the full antenna grid of 121 antennas. The splitting process is based on the energy of the particles, where the amount of particles per part can vary from $540,000$ (energies below $10^{7}$ eV) down to only 1 (starting at energies of $10^{13}$ eV). Figure 18 shows the distribution of the CPU time of the $591$ cores used for the in-ice simulation. Most of the CPUs manage to stay within $10$ h of CPU time. Some cores however take significantly longer, driving up the job run time to approximately $24$ h. The distribution shows there is still room for optimization in the splitting process. Combining the particle groups corresponding to the simulations with low CPU time would decrease the amount of cores needed, while splitting up the particle groups corresponding to the simulations with high CPU time would decrease the job run time. The longest CPU time for a simulation of the propagation of a single particle was $12.2$ h, which means for this specific case the job run time will always be at least $12.2$ h, unless the antenna grid is distributed over different cores as well. The total CPU time was $2985$ h.